# Assessing treatment efficacy for interval-censored endpoints using multistate semi-Markov models fit to multiple data streams

### v2

@@ 2.1 Motivating example: REGEN-COV mAb for COVID-19 prophylaxis
Table 1(b) summarizes the assumptions under which we formulate a biological model for infection, symptoms, and immunological response. Evidence of SARS-CoV-2 infection may be captured by positive PCR, manifestation of symptoms, or positive serology. Participants who never manifest symptoms or test positive are assumed to be infection naïve at the end of the EAP, i.e., unaffected implies never infected. In this spirit, we treat infection as the occurrence of a participant being measurably affected, as evidenced by viral shedding detectable by nasopharyngeal RT-qPCR. Thus, in our model, infection is the transition from from naïve to PCR+. The other modeling assumptions reflect our understanding of the temporal precedence of viral shedding, symptoms, and seroconversion. For instance, SARS-CoV-2 infected individuals shed detectable virus prior to manifesting symptoms (Puhach and others, 2023). These assumptions inform the formulation of the model states and allowable transitions depicted in Figure 1(c), and identify the model state, or set of possible model states, at each observation time, as shown in the last two columns of Table 1(a).
@@ 2.2 Data for continuously and intermittently observed multistate processes
The data for participant $i$ accrue at times $t_{i,0},\dots,t_{i,L_{i}}$. Let $O_{i,j}=0$ if we observe $J(t_{i,j})\in\mathcal{S}(t_{i,j})$, where $\mathcal{S}(t_{i,j})\subseteq\mathcal{S}$ are states consistent with the data at $t_{i,j}$. Let $O_{i,j}=1$ if the $Z_{i}$ is continuously observed over $(t_{i,{j-1}},t_{i,j}]$. Specifically we observe $J(\tau)\in\mathcal{S}(\tau)$ for $\tau\in(t_{i,j-1},t_{i,j}]$, which we denote using the shorthand, $J(t_{i,j-1},t_{i,j})\in\mathcal{S}(t_{i,j-1},t_{i,j})$. In Figure 2(b), we observe that the participant is healthy from $t_{0}$ to $T_{1}$, when their disease recurs, and that they are ill during $(T_{1},t_{1}]$. We then observe that the participant is alive at $t_{2}$, but not whether their disease has recurred, so $\mathcal{S}(t_{2})=\{\mathrm{Healthy},\mathrm{Ill}\}$. The participant is known to have died by $t_{3}$. Let $\mathbf{Y}_{i,j}=\left(\mathcal{S}(t_{i,j-1},t_{i,j}),O_{i,j}\right)$ denote the data for participant $i$ over the interval $(t_{i,j-1},t_{i,j}]$, and $\mathbf{Y}_{i}=\left(\mathbf{Y}_{i,j}\right)_{j=0,\dots,L_{i}}$. Finally, $\mathbf{Y}=\left(\mathbf{Y}_{i}\right)_{i=1,\dots,N}$ is the data for all study participants.
@@ 2.4 Maximum likelihood via Monte Carlo expectation-maximization
Direct optimization of the marginal likelihood is challenging when data are intermittently observed since the number and timing of transitions are unknown. The expectation-maximization (EM) algorithm of Dempster and others (1977) facilitates maximum likelihood estimation in missing data settings by alternating between an E-step that calculates the expected complete data log-likelihood, referred to as the Q-function, and an M-step that maximizes the Q-function. Let $\widehat{\boldsymbol{\theta}}^{(t-1)}$ be the estimate of $\boldsymbol{\theta}$ after $t-1$ steps of the EM algorithm. The Q-function is,
$Q\left(\boldsymbol{\theta},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$ $=\mathrm{E}_{\mathbf{Z}\mid\mathbf{Y},\mathbf{X},\widehat{\boldsymbol{\theta}}^{(t-1)}}\left[\ell\left(\mathbf{Z},\mathbf{Y}\mid\mathbf{X},\boldsymbol{\theta}\right)\right],$ (7)
@@
$\widetilde{Q}\left(\boldsymbol{\theta},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$ $=\sum_{i=1}^{P}\frac{\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t-1)}l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}\right)}{\sum_{j=1}^{M_{i}}\nu_{i,j}^{(t-1)}},$ (8)
@@
$\nu_{i,m}^{(t-1)}=\frac{f_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\boldsymbol{\theta}^{\prime}\right)},$ (9)
@@
$l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}\right)$ $=\log\left(f_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)p_{i}(\mathbf{Y}_{i}\mid Z_{i}^{(m)},\mathbf{X}_{i},\boldsymbol{\theta})\right),$
$=\log\left(f_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)\mathbb{I}\left\{Z_{i}^{(m)}\textsf{ concords with }\mathbf{Y}_{i}\right\}\right),$
$=\log\left(f_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)\right),$
@@
The ascent property of the EM algorithm that assures $L\left(\mathbf{Y}\mid\widehat{\boldsymbol{\theta}}^{(t)}\right)\geq L\left(\mathbf{Y}\mid\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$ is no longer guaranteed in MCEM. We implement the ascent MCEM algorithm of Caffo and others (2005), which recovers the ascent property with high probability by accepting the maximizer in each M-step only when the increase in the Q-function is distinguished from Monte Carlo error with sufficient confidence. The ascent MCEM algorithm calls to augment the pool of Monte Carlo samples when the change in the Q-function is statistically indistinguishable from zero. The Monte Carlo standard of the change in the Q-function is a function of the effective sample size (ESS),
@@
Once the MCEM algorithm has converged, uncertainty about the MLE, $\widehat{\boldsymbol{\theta}}$, can be quantified using bootstrap confidence intervals or asymptotic model-based intervals, which can be obtained by inverting the observed Fisher information matrix. In the latter case, expectations of the score and Hessian required to compute the observed Fisher information can be estimated using recycling sample paths generated in the MCEM algorithm (see Section A of the Supplementary Material). Mandel (2013) details a convenient asymptotic Monte Carlo procedure, similar to a Bayesian posterior predictive procedure, which quantifies uncertainty for functionals of interest, such as expected sojourn times, by sampling $\widehat{\boldsymbol{\theta}}^{\star}$ from the asymptotic distribution of $\widehat{\boldsymbol{\theta}}$ and then simulating sample paths using $\widehat{\boldsymbol{\theta}}^{\star}$, which are then summarized. Monte Carlo confidence intervals are obtained by calculating quantiles of the summaries of interest.
@@ 2.4.1 Data-conditioned Markov proposals for semi-Markov sample paths
Let $\{\widetilde{Z}(\tau),\tau\geq 0\}$ be a continuous-time Markov process taking values in $\mathcal{S}$, and let $\widetilde{\boldsymbol{\theta}}=\left(\widetilde{\boldsymbol{\gamma}},\widetilde{\boldsymbol{\beta}}\right)$ be a vector of Markov model parameters that we again partition into those pertaining to the baseline transition intensities and covariates. A proportional transition intensity parameterization with a time-homogeneous baseline intensity is convenient. In contrast to (2), the transition intensity is now independent of the sojourn time, $\tau_{n}$. Hence,
$\lambda_{k\ell}\left(\tau,\mathcal{H}(\tau^{-}),\mathbf{X}(\tau^{-}),\widetilde{\boldsymbol{\theta}}\right)=\widetilde{\gamma}_{k\ell}\exp\left(\mathbf{X}(\tau^{-})\boldsymbol{\widetilde{\boldsymbol{\beta}}}_{k\ell}\right).$ (11)
Suppressing the dependence on $\mathcal{H}(\tau^{-})$, $\mathbf{X}$ and $\widetilde{\boldsymbol{\theta}}$, let $\boldsymbol{\Lambda}(\tau)=(\lambda_{k\ell}(\tau))_{k,\ell\in\mathcal{S},}$, denote the transition intensity matrix under a Markov model, with $\lambda_{kk}(\tau)=-\sum_{\ell\neq k}\lambda_{k\ell}(\tau)$. The transition probability matrix, $\mathbf{P}(\tau_{0},\tau_{1})=(p_{k\ell}(\tau_{0},\tau_{1}))_{k,\ell\in\mathcal{S}}$, where $p_{k\ell}(\tau_{0},\tau_{1})=\Pr(Z(\tau_{1})=\ell\mid Z(\tau_{0})=k)$, solves the Kolmogorov forward equation (KFE),
@@
The likelihood for a Markov model can be maximized using standard optimization routines to yield a maximum likelihood estimate (MLE) of $\widetilde{\boldsymbol{\theta}}$, under which we propose sample paths.
@@
            THEN append observed path: $\widetilde{Z}\Leftarrow(\widetilde{Z},Y_{i}(t_{i,j-1},t_{i,j}))$.
@@
            THEN sample $\widetilde{Z}(t_{i,j-1},t_{i,j})\mid J(t_{i,j-1}),J(t_{i,j})$ via uniformization
            AND append sampled path: $\widetilde{Z}\Leftarrow(\widetilde{Z},\widetilde{Z}(t_{i,j-1},t_{i,j}))$.
Return path: $\widetilde{Z}_{i}=(J,W)_{i}$, over $[t_{i,0},t_{i,L_{i}}]$.
@@ 2.5 Model comparison and marginal likelihood
$\widehat{g}\left(\mathbf{Y}\mid\mathbf{X},\boldsymbol{\theta}^{(t)}\right)=\prod_{i}^{n}r_{i}(\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}^{\prime})\frac{1}{M_{i}}\sum_{m=1}^{M_{i}}\nu^{(t)}_{i,m},$ (14)
@@ 3.1 Analysis of intercurrent events in an illness-death model
The progressive illness-death model is useful in the study of diseases, such as cancer, where death competes with disease recurrence. Patients begin follow-up when they are deemed healthy following successful treatment. The model has three states – disease free (healthy), disease recurrence (ill), and dead – and three transitions –- healthy to ill, healthy to dead, and ill to dead (Figure S2). Disease recurrence is interval-censored but the time of death and its cause are observed.
We simulated clinical histories for 250 participants in 1,000 hypothetical studies using an illness-death model with Weibull transition intensities with shape parameter greater than 1, so each intensity was 0 at the time of state entry and increased over time (Table S5). Clinical data was accrued at baseline, which was the time of entry into the healthy state, and at random times spaced roughly every 1 or 3 months over 1 year of follow-up. Confidence intervals for all quantities of interest were computed by summarizing sample paths simulated from the asymptotic distribution of the MLEs. Additional details of the simulation setup and models fit are given in Appendix B.
@@ 3.2 Estimating protective efficacy and natural history in simulated trials
The analysis of the REGEN-2069 trial data in Section 4 aims to characterize the effect of mAb on infection and seroconversion, and to quantify the duration of PCR positivity. Inference is based on the model, depicted in Figure 1(c), which has a naïve uninfected state, and four post-infection states characterized by symptom history and whether SARS-CoV-2 infection is, or has ceased to be, detectable by PCR. A two-stage procedure for calculating the probability of seroconversion given infection is detailed in Section 4.2.1. The data consist of symptom onset times, seroconversion at four weeks, and PCR measurements roughly every week throughout the efficacy assessment period (EAP), with some variation in the visit times to emulate a realistic trial. Trials were simulated from a more complicated semi-Markov model with nine states – a naïve uninfected state, and eight post-infection states characterized by symptom history, PCR positivity, and seroconversion (Table S8). The transitions from naïve to PCR+ without symptoms and from PCR+ without symptoms to diseased were assigned Weibull transition intensities, while the remainder of the transitions had exponential intensities. The shape parameter for the infection transition, $1\rightarrow 2$, was less than 1 to reflect a decreasing force of infection, and the shape parameter for symptom onset was greater than 1 to induce a delay in symptom onset. We set the simulation parameters so that subjects receiving mAb have lower rates of infection and seroconversion along with shorter durations of PCR positivity. The parameterization of the simulation model and ground truths for key quantities of interest are summarized in Tables S9 and S10.
We assesed our ability to recover key functionals with two formulations of the five state inferential model: the first had Weibull baseline intensities for transitions from naïve to PCR+ without symptoms ($1\rightarrow 2$) and from PCR+ without symptoms to diseased ($2\rightarrow 4$) and with exponential intensities from PCR+ to PCR-- states ($2\rightarrow 3$ and $4\rightarrow 5$), or another with linear spline baseline intensities with an interior knot at five days for the $1\rightarrow 2$ transition and an interior knot at one week for all other transitions (Table S11). As benchmarks, we report crude estimates obtained by naïvely tabulating events in the data as would be typically done in a descriptive analysis (Appendix C.3), and we also fit the 9 state simulation model to a more granular version of data where the state at each PCR assessment was known. This model is not fully identifiable with real-world data where serology is only assessed at day 28 since the order of seroconversion and cessation of detectability by PCR is unknown. Confidence intervals for all quantities were obtained via the Bayesian bootstrap (Rubin, 1981) and the bootstrap weights were stratified by treatment.
@@
Table 3: Evaluation of estimates of key quantities involving infection in 1,000 simulated trials. Relative bias and confidence interval width are scaled by the true value. The simulation model is the 9 state generative model described in Supplementary Appendix C.1. The parameterization of transition intensities in the 5 state models is given in Table S11. The collapsed parametric model has Weibull baseline intensities for infection and symptom onset, and exponential intensities for exiting states of detectability by PCR. The collapsed semi-parametric model has degree 1 B-splines for all baseline intensities. Crude estimate refers to simple tabulation of observed quantities as described in Section C.3. The simulation model is fit to hypothetical panel data where serology is assessed along with PCR at weekly intervals, whereas the collapsed models are fit to data that follows the design of REGEN-2069 where serology was assessed at day 28.
@@
Crude estimates of quantities implicated by asymptomatic infection were biased and had unacceptably low coverage. The 9 state simulation model, unsurprisingly, yielded unbiased estimates and 95% CIs that attained their nominal coverage. Estimates obtained with both of the 5-state models had negligible bias and good coverage for PE against infection, including asymptomatic infection. There was about 6% relative bias in the PE for asymptomatic infection, which is attributable to slight underestimation of extent of asymptomatic infection on the placebo arm (Tables S13 and S14). The relative risk (RR) of symptoms following infection and RR of seroconversion following infection were well estimated. Both models exhibited slight bias for the restricted mean time to infection and duration of PCR positivity along with modest under-coverage for these quantities and the probability of being detected by PCR. This is attributable to some individuals who are seropositive at the end of the EAP but who never develop symptoms or test PCR+. The reduced model implicitly treats the timing of their infection as random from the distribution of infection times. This results in a slight bias that favors early infection under the fitted model, which is exacerbated in the placebo arm by virtue of its higher incidence of infections that are only detected by serology. Overall, the two stage estimation procedure using the reduced 5 state model has good operating characteristics for estimating key quantities of interest.

### v3

@@ 2.1 Motivating example: REGEN-COV mAb for COVID-19 prophylaxis
Table 1(b) summarizes the assumptions under which we formulate a biological model for infection, symptoms, and immunological response. Evidence of SARS-CoV-2 infection may be captured by positive PCR, manifestation of symptoms, or positive serology. Participants who never manifest symptoms or test positive are assumed to be infection naïve at the end of the EAP, i.e., unaffected implies never infected. In this spirit, we treat infection as the occurrence of a participant being measurably affected, as evidenced by viral shedding detectable by nasopharyngeal RT-qPCR. Thus, in our model, infection is the transition from naïve to PCR+. The other modeling assumptions reflect our understanding of the temporal precedence of viral shedding, symptoms, and seroconversion. For instance, SARS-CoV-2 infected individuals shed detectable virus prior to manifesting symptoms (Puhach and others, 2023). These assumptions inform the formulation of the model states and allowable transitions depicted in Figure 1(c), and identify the model state, or set of possible model states, at each observation time, as shown in the last two columns of Table 1(a).
@@ 2.2 Data for continuously and intermittently observed multistate processes
The data for participant $i$ accrue at times $t_{i,0},\dots,t_{i,L_{i}}$. Let $O_{i,j}=0$ if we observe $J(t_{i,j})\in\mathcal{S}(t_{i,j})$, where $\mathcal{S}(t_{i,j})\subseteq\mathcal{S}$ are states consistent with the data at $t_{i,j}$. Let $O_{i,j}=1$ if the $Z_{i}$ is continuously observed over $(t_{i,{j-1}},t_{i,j}]$. Specifically, we observe $J(\tau)\in\mathcal{S}(\tau)$ for $\tau\in(t_{i,j-1},t_{i,j}]$, which we denote using the shorthand, $J(t_{i,j-1},t_{i,j})\in\mathcal{S}(t_{i,j-1},t_{i,j})$. In Figure 2(b), we observe that the participant is healthy from $t_{0}$ to $T_{1}$, when their disease recurs, and that they are ill during $(T_{1},t_{1}]$. We then observe that the participant is alive at $t_{2}$, but not whether their disease has recurred, so $\mathcal{S}(t_{2})=\{\mathrm{Healthy},\mathrm{Ill}\}$. The participant is known to have died by $t_{3}$. Let $\mathbf{Y}_{i,j}=\left(\mathcal{S}(t_{i,j-1},t_{i,j}),O_{i,j}\right)$ denote the data for participant $i$ over the interval $(t_{i,j-1},t_{i,j}]$, and $\mathbf{Y}_{i}=\left(\mathbf{Y}_{i,j}\right)_{j=0,\dots,L_{i}}$. Finally, $\mathbf{Y}=\left(\mathbf{Y}_{i}\right)_{i=1,\dots,N}$ is the data for all study participants.
@@ 2.4 Maximum likelihood via Monte Carlo expectation-maximization
Direct optimization of the marginal likelihood is challenging when data are intermittently observed since the number and timing of transitions are unknown. The expectation-maximization (EM) algorithm of Dempster and others (1977) facilitates maximum likelihood estimation in missing data settings by alternating between an E-step that calculates the expected complete data log-likelihood, referred to as the Q-function, and an M-step that maximizes the Q-function. Let $\hat{\boldsymbol{\theta}}^{(t-1)}$ be the estimate of $\boldsymbol{\theta}$ after $t-1$ steps of the EM algorithm. The Q-function is,
$Q\left(\boldsymbol{\theta},\hat{\boldsymbol{\theta}}^{(t-1)}\right)$ $=\mathrm{E}_{\mathbf{Z}\mid\mathbf{Y},\mathbf{X},\hat{\boldsymbol{\theta}}^{(t-1)}}\left[\ell\left(\mathbf{Z},\mathbf{Y}\mid\mathbf{X},\boldsymbol{\theta}\right)\right],$ (7)
@@
$\tilde{Q}\left(\boldsymbol{\theta},\hat{\boldsymbol{\theta}}^{(t-1)}\right)$ $=\sum_{i=1}^{P}\frac{\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t-1)}l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}\right)}{\sum_{j=1}^{M_{i}}\nu_{i,j}^{(t-1)}},$ (8)
@@
$\nu_{i,m}^{(t-1)}=\frac{f_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\boldsymbol{\theta}^{\prime}\right)},$ (9)
@@
$l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}\right)$ $=\log\left(f_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)p_{i}(\mathbf{Y}_{i}\mid Z_{i}^{(m)},\mathbf{X}_{i},\boldsymbol{\theta})\right),$
$=\log\left(f_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)\mathbb{I}\left\{Z_{i}^{(m)}\textsf{ concords with }\mathbf{Y}_{i}\right\}\right),$
$=\log\left(f_{i}\left(Z_{i}^{(m)}\mid\mathbf{X}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)\right),$
@@
The ascent property of the EM algorithm that assures $L\left(\mathbf{Y}\mid\hat{\boldsymbol{\theta}}^{(t)}\right)\geq L\left(\mathbf{Y}\mid\hat{\boldsymbol{\theta}}^{(t-1)}\right)$ is no longer guaranteed in MCEM. We implement the ascent MCEM algorithm of Caffo and others (2005), which recovers the ascent property with high probability by accepting the maximizer in each M-step only when the increase in the Q-function is distinguished from Monte Carlo error with sufficient confidence. The ascent MCEM algorithm calls to augment the pool of Monte Carlo samples when the change in the Q-function is statistically indistinguishable from zero. The Monte Carlo standard of the change in the Q-function is a function of the effective sample size (ESS),
@@
Once the MCEM algorithm has converged, uncertainty about the MLE, $\hat{\boldsymbol{\theta}}$, can be quantified using bootstrap confidence intervals or asymptotic model-based intervals, which can be obtained by inverting the observed Fisher information matrix. In the latter case, expectations of the score and Hessian required to compute the observed Fisher information can be estimated using recycling sample paths generated in the MCEM algorithm (see Section A of the Supplementary Material). Mandel (2013) details a convenient asymptotic Monte Carlo procedure, similar to a Bayesian posterior predictive procedure, which quantifies uncertainty for functionals of interest, such as expected sojourn times, by sampling $\hat{\boldsymbol{\theta}}^{\star}$ from the asymptotic distribution of $\hat{\boldsymbol{\theta}}$ and then simulating sample paths using $\hat{\boldsymbol{\theta}}^{\star}$, which are then summarized. Monte Carlo confidence intervals are obtained by calculating quantiles of the summaries of interest.
@@ 2.4.1 Data-conditioned Markov proposals for semi-Markov sample paths
Let $\{\tilde{Z}(\tau),\tau\geq 0\}$ be a continuous-time Markov process taking values in $\mathcal{S}$, and let $\tilde{\boldsymbol{\theta}}=\left(\tilde{\boldsymbol{\gamma}},\tilde{\boldsymbol{\beta}}\right)$ be a vector of Markov model parameters that we again partition into those pertaining to the baseline transition intensities and covariates. A proportional transition intensity parameterization with a time-homogeneous baseline intensity is convenient. In contrast to (2), the transition intensity is now independent of the sojourn time, $\tau_{n}$. Hence,
$\lambda_{k\ell}\left(\tau,\mathcal{H}(\tau^{-}),\mathbf{X}(\tau^{-}),\tilde{\boldsymbol{\theta}}\right)=\tilde{\gamma}_{k\ell}\exp\left(\mathbf{X}(\tau^{-})\boldsymbol{\tilde{\boldsymbol{\beta}}}_{k\ell}\right).$ (11)
Suppressing the dependence on $\mathcal{H}(\tau^{-})$, $\mathbf{X}$ and $\tilde{\boldsymbol{\theta}}$, let $\boldsymbol{\Lambda}(\tau)=(\lambda_{k\ell}(\tau))_{k,\ell\in\mathcal{S},}$, denote the transition intensity matrix under a Markov model, with $\lambda_{kk}(\tau)=-\sum_{\ell\neq k}\lambda_{k\ell}(\tau)$. The transition probability matrix, $\mathbf{P}(\tau_{0},\tau_{1})=(p_{k\ell}(\tau_{0},\tau_{1}))_{k,\ell\in\mathcal{S}}$, where $p_{k\ell}(\tau_{0},\tau_{1})=\Pr(Z(\tau_{1})=\ell\mid Z(\tau_{0})=k)$, solves the Kolmogorov forward equation (KFE),
@@
The likelihood for a Markov model can be maximized using standard optimization routines to yield a maximum likelihood estimate (MLE) of $\tilde{\boldsymbol{\theta}}$, under which we propose sample paths.
@@
            THEN append observed path: $\tilde{Z}\Leftarrow(\tilde{Z},Y_{i}(t_{i,j-1},t_{i,j}))$.
@@
            THEN sample $\tilde{Z}(t_{i,j-1},t_{i,j})\mid J(t_{i,j-1}),J(t_{i,j})$ via uniformization
            AND append sampled path: $\tilde{Z}\Leftarrow(\tilde{Z},\tilde{Z}(t_{i,j-1},t_{i,j}))$.
Return path: $\tilde{Z}_{i}=(J,W)_{i}$, over $[t_{i,0},t_{i,L_{i}}]$.
@@ 2.5 Model comparison and marginal likelihood
$\hat{g}\left(\mathbf{Y}\mid\mathbf{X},\boldsymbol{\theta}^{(t)}\right)=\prod_{i}^{n}r_{i}(\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}^{\prime})\frac{1}{M_{i}}\sum_{m=1}^{M_{i}}\nu^{(t)}_{i,m},$ (14)
@@ 3.1 Analysis of intercurrent events in an illness-death model
The progressive illness-death model is useful in the study of diseases, such as cancer, where death competes with disease recurrence. Patients begin follow-up when they are deemed healthy following successful treatment. The model has three states – disease free (healthy), disease recurrence (ill), and dead – and three transitions –- healthy to ill, healthy to dead, and ill to dead (Figure S2). Disease recurrence is interval-censored, but the time of death and its cause are observed.
We simulated clinical histories for 250 participants in 1,000 hypothetical studies using an illness-death model with Weibull transition intensities with shape parameter greater than 1, so each intensity was 0 at the time of state entry and increased over time (Tables S1 and S2). Clinical data was accrued at baseline, which was the time of entry into the healthy state, and at random times spaced roughly every 1 or 3 months over 1 year of follow-up. The data was analyzed using a Markov model with exponential transition intensities and three semi-Markov models, which included a model with Weibull intensities and two models with B-spline intensities (see Table S3 for parameterizations). Confidence intervals for all quantities of interest were computed by summarizing sample paths simulated from the asymptotic distribution of the MLEs. Additional details of the simulation setup and models fit are given in Appendix B.
@@ 3.2 Estimating protective efficacy and natural history in simulated trials
The analysis of the REGEN-2069 trial data in Section 4 aims to characterize the effect of mAb on infection and seroconversion, and to quantify the duration of PCR positivity. Inference is based on the model, depicted in Figure 1(c), which has a naïve uninfected state, and four post-infection states characterized by symptom history and whether SARS-CoV-2 infection is, or has ceased to be, detectable by PCR. A two-stage procedure for calculating the probability of seroconversion given infection is detailed in Section 4.2.1. The data consist of symptom onset times, seroconversion at four weeks, and PCR measurements roughly every week throughout the efficacy assessment period (EAP), with some variation in the visit times to emulate a realistic trial. Trials were simulated from a more complicated semi-Markov model with nine states – a naïve uninfected state, and eight post-infection states characterized by symptom history, PCR positivity, and seroconversion (Table S4). The transitions from naïve to PCR+ without symptoms and from PCR+ without symptoms to diseased were assigned Weibull transition intensities, while the remainder of the transitions had exponential intensities. The shape parameter for the infection transition, $1\rightarrow 2$, was less than 1 to reflect a decreasing force of infection, and the shape parameter for symptom onset was greater than 1 to induce a delay in symptom onset. We set the simulation parameters so that subjects receiving mAb have lower rates of infection and seroconversion along with shorter durations of PCR positivity. The parameterization of the simulation model and ground truths for key quantities of interest are summarized in Tables S5 and S6.
We assessed our ability to recover key functionals with two formulations of the five state inferential model: the first had Weibull baseline intensities for transitions from naïve to PCR+ without symptoms ($1\rightarrow 2$) and from PCR+ without symptoms to diseased ($2\rightarrow 4$) and with exponential intensities from PCR+ to PCR-- states ($2\rightarrow 3$ and $4\rightarrow 5$), or another with linear spline baseline intensities with an interior knot at five days for the $1\rightarrow 2$ transition and an interior knot at one week for all other transitions (Table S7). As benchmarks, we report crude estimates obtained by naïvely tabulating events in the data as would be typically done in a descriptive analysis (Appendix C.3), and we also fit the 9 state simulation model to a more granular version of data where the state at each PCR assessment was known. This model is not fully identifiable with real-world data where serology is only assessed at day 28 since the order of seroconversion and cessation of detectability by PCR is unknown. Confidence intervals for all quantities were obtained via the Bayesian bootstrap (Rubin, 1981) and the bootstrap weights were stratified by treatment. Simulation results are reported in Tables 2, S8, S9, S10, and S11.
@@
Table 2: Evaluation of estimates of key quantities involving infection in 1,000 simulated trials. Relative bias and confidence interval width are scaled by the true value. The simulation model is described in Supplementary Appendix C.1. The parameterization of transition intensities in the 5 state models is given in Table S7. The collapsed parametric model has Weibull baseline intensities for infection and symptom onset, and exponential intensities for exiting states of detectability by PCR. The collapsed semi-parametric model has degree 1 B-splines for all baseline intensities. The simulation model was fit to data where serology was assessed weekly along with PCR, whereas the collapsed models and crude estimates used data that followed the design of REGEN-2069 where serology was assessed at day 28.
@@
Crude estimates of quantities implicated by asymptomatic infection were biased and had unacceptably low coverage. The 9 state simulation model, unsurprisingly, yielded unbiased estimates and 95% CIs that attained their nominal coverage. Estimates obtained with both of the 5-state models had negligible bias and good coverage for PE against infection, including asymptomatic infection. There was about 6% relative bias in the PE for asymptomatic infection, which is attributable to slight underestimation of extent of asymptomatic infection on the placebo arm (Tables S9 and S10). The relative risk (RR) of symptoms following infection and RR of seroconversion following infection were well estimated. Both models exhibited slight bias for the restricted mean time to infection and duration of PCR positivity along with modest under-coverage for these quantities and the probability of being detected by PCR. This is attributable to some individuals who are seropositive at the end of the EAP, but who never develop symptoms or test PCR+. The reduced model implicitly treats the timing of their infection as random from the distribution of infection times. This results in a slight bias that favors early infection under the fitted model, which is exacerbated in the placebo arm by virtue of its higher incidence of infections that are only detected by serology. Overall, the two stage estimation procedure using the reduced 5 state model has good operating characteristics for estimating key quantities of interest.
3.3 Comparison with existing approaches
In Section E.1 of the Supplementary Material, we compare our approach with that taken in Aralis and Brookmeyer (2019), who generate paths from a semi-Markov model by forward simulation and reject proposed paths that are inconsistent with the data. We explore the performance of the two approaches for fitting a progressive illness-death model and find that data-conditioned path proposals yield dramatic efficiency gains as the number of observations per-individual increases (Table S16). The relative efficiency gains of our approach are greater for more complex models since the rejection sampler becomes increasingly likely to propose paths that are inconsistent with the data.
In Section E.2 of the Supplementary Material, we compare semi-parametric models fit using our approach with phase-type models that are commonly used to approximate semi-Markov models in panel data settings. This latter class of models uses a latent Coxian phase-type structure to approximate arbitrary time-to-event distributions and is amenable to numerical optimization (Titman and Sharples, 2010). Phase-type models can struggle to approximate intensities that are initially close to zero and remedies for ameliorating the ensuing lack of fit introduce additional technical complications and computational challenges. We demonstrate that approximating the transition intensity for illness-onset in a semi-Markov illness-death process using a B-spline with a single interior knot yields a substantially improved fit compared with a phase-type model with two latent states to capture the semi-Markov dynamics of this transition (Table S17). In particular, expected prevalence of illness is too early and too low under the phase-type model, whereas the semi-Markov model accurately tracks the true illness prevalence over time (Figure S5). The phase-type model still vastly outperforms a simple Markov model and we speculate that phase-type models could be used as an alternative proposal distribution within our MCEM framework in settings where simple time-homogeneous Markov models are inefficient. Phase-type models are characterized by a latent Markovian structure that, like simpler Markov dynamics, lends itself to data-conditioned simulation. Whether a more complex alternative proposal distribution is warranted could be determined based on having a small ESS relative to the nominal number of paths.